# Triple-BERT: Do We Really Need MARL for Order Dispatch on Ride-Sharing Platforms?

### v1

@@ 3.3.2 Stage 2: Centralized TD3 Training
(i) Actor: To address the large action space, we propose an action decomposition method that separates the probability of selecting each worker-order assignment combination into the probabilities of each worker choosing their respective orders. First, we expand the utility matrix $M_{t}$ output by the Actor QK-Attention to $\mathcal{M}_{t}=[M_{t},N_{t}]\in\mathbb{R}^{n,m_{t}+1}$, where $N_{t}$ is an $n$-dimensional vector representing the utility of each worker choosing no order. This vector can be obtained by processing the output of Actor-BERT with a MLP, i.e., $N_{t}=\text{MLP}([\overline{w}_{1,t},\overline{w}_{2,t},\ldots,\overline{w}_{n,t}])$. This allows us to compute the probability of each worker choosing each action using a logit model [19], if the actions among workers are independent, i.e., $\mathscr{P}_{t}=\text{Softmax}(\mathcal{M}_{t})$. According to this independent assumption, the joint action probability can be expressed as: $\prod_{i,j\in\text{h}(A_{t})}\mathscr{P}_{i,j,t}$, where $\text{h}()$ is defined as the function $\text{h}(A_{t})=\{(i,j)|a_{i,j,t}=1\}$. Similar to stage 1, we set the probability of those unavailable workers choosing no order to 1 and all other actions to 0 in $\mathscr{P}_{t}$. (We consider this independent assumption can be approximately realized after the network is well-trained, as BERT has already captured the relationships among workers, including their strategic interactions.)
However, since an order cannot be assigned to different workers repeatedly, the actions among workers are actually not independent. Intuitively, if a worker is more willing to choose a particular action, this action should have a higher probability of being selected by this worker in the joint action. Based on this intuition, the action choosing probability can be defined as:
@@
where $\text{z}(\cdot)$ is an increasing function that also depends on the current state $S_{t}$ (which we omit for simplicity). This equation implies that if an action $A_{t}$ has a higher value of $\prod_{i,j\in\text{h}(A_{t})}\mathscr{P}_{i,j,t}$, it will have a higher probability of being chosen.
@@
Optimizing this probability using vanilla TD3 is challenging due to the variable action space and the gap between action probabilities and the selected action (the gradient cannot propagate through them). To address this, we employ an approximate policy gradient optimization method [26]:
@@ 4 Experiment
Method DeepPool [1] BMG-Q [16] HIVES [12] Enders et al. [8] CEVD [3] Triple-BERT Type Independent CTDE Centralized RL Algorithm IDDQN [28] IDDQN [28] QMIX [23] MASAC [11] VD [^2] The original VD is a CTDE method. However, the CEVD variant modifies it to a centralized version. [25] TD3 [10] Multi-Agent $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ Network Backbone MLP [^3] In the original paper fo DeepPool, the authors used CNN. However, due to differences in the observation space of our task, we replaced it with MLP. GAT [29] GRU [5] MLP+Attention MLP BERT [7] Model Size 20K 117K 16M 118K 23K 16M GPU Occupation (GB) 3.97 4.28 6.01 8.19 21.45 8.03 Average Reward ($10^{3}$) 12.72 13.04 12.37 12.04 13.16 14.73
@@
To validate the proposed method, we evaluate its performance in the ride sharing dispatching task using real-world yellow ride-hailing data from Manhattan, New York City[27] [^4] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page . To illustrate the efficiency and superiority of our proposed Triple-BERT, we compare it with several previous ride sharing methods of different types, including Independent MARL, CTDE MARL, and Centralized MARL, as shown in Table 1. Detailed information regarding our experiment configuration, simulator setup, and a comprehensive description of the comparative experiment can be found in Appendix E.
@@
Figure 3: Training Process: Each method is trained three times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$. The shaded area represents the range of fluctuations, while the solid line indicates the average value.
@@
As shown in Fig. 3, we first illustrate the training process of different models by evaluating their performance in the training scenario every 10 episodes. The six sub-figures depict the cumulative reward, the number of orders served, and the average delivery time, detour time, pickup time, and confirmation time for each order. Additionally, the Greedy method serves as a baseline, where orders are assigned to the nearest worker. It is evident that our method outperforms the other models in most metrics, with the cumulative reward exceeding that of the best alternative method by approximately 15%. The highest number of served orders indicates that our method achieves better cooperation among workers. We then evaluate these methods over different periods, and the average rewards are shown in Table 1, where our method also demonstrates the best performance. More details about the experimental results can be found in Appendix E.4.
@@
To further demonstrate the model’s efficiency, we conduct a series of ablation studies. In terms of model training, we compare the performance of the model with and without stage 1 pre-training. Regarding the network structure, we primarily compare the QK-Attention mechanism with and without the proposed positive normalization module. The detailed results are shown in Fig. 3. We observe that without stage 1 pre-training, the model fails to converge and exhibits significant fluctuations. Particularly in the later stages, the reward begins to decrease, which can be attributed to the lack of samples. Additionally, without the proposed normalization in QK-Attention, the model performs poorly, underperforming compared to all other methods. This is due to parameter redundancy, which leads to substantial fluctuations and hinders efficient learning.
@@ Appendix C Policy Gradient Proof
According to the policy gradient theory [26], we have:
@@
Since $\text{z}(x)$ is an increasing function, the elasticity is always non-negative. Here, we assume that the elasticity of $\text{z}(x)$ with respect to $x$ can be approximately viewed as a positive constant, which means that $\text{z}(x)$ has the format $ax^{b}$ (where $a,b>0$). Thus, we have: $\nabla_{\Theta}\text{J}(\Theta)\propto\mathbb{E}_{\pi_{\Theta}^{T}}\left[\left(\text{Q}^{TD3}_{\pi_{\Theta}^{T}}(S_{t},A_{t})-B\right)\nabla_{\Theta}\sum_{i,j\in\text{h}(A_{t})}\log\mathscr{P}_{i,j,t}\right]$, corresponding to Eq. 7.
Actually, the independent action probability does not exist in a real sense. However, for any given definition of ’z’, a corresponding independent action probability can be established. We believe this is a valid design since ’z’ is an increasing function. Intuitively, if a driver-order pair has a higher probability of being chosen in the independent condition, it should have a higher utility and thus a higher probability of being selected in the joint action. This ’z’ function possesses the advantageous property of maintaining a positive constant elasticity, allowing us to eliminate the elasticity term $\epsilon$ in Eq. 14 without affecting the gradient direction. As a result, when training our method using the gradient in Eq. 7, the definition of the ’z’ function is implicit as given in $ax^{b}$ (where $a,b>0$). Given our method’s SOTA performance, we consider this approach to be effective.
@@
As mentioned in Section 3.3.2, during training, we add random noise to $\mathscr{P}_{t}$ and then choose the action that maximizes $\sum_{i,j\in\text{h}(A_{t})}\log\mathscr{P}_{i,j,t}$. Currently, the mapping from $\sum_{i,j\in\text{h}(A_{t})}\log\mathscr{P}_{i,j,t}$ to the choosing probability $\pi_{\Theta}^{T}$ corresponds to $\text{z}(\cdot)$. To further illustrate the robustness of our method, we compare the performance of our model using Gaussian noise, uniform noise, and binary symmetric channel (BSC) noise, where the noise follows a Bernoulli distribution and has been widely utilized in previous work [16; 15]. During training, we gradually reduce the noise to make the policy more deterministic. The experimental results are shown in Fig. 4, where we observe that, despite certain performance differences between the various types of noise, they all outperform conventional MARL methods. This suggests the efficiency and high robustness of our proposed method, indicating that the detailed expression of $\text{z}(x)$ does not significantly influence the validation of the method based on Eq. 7, even if it may cause some performance gaps. The optimal noise for our task may require further exploration. For fairness, we choose to use BSC noise when comparing with other methods, even though it appears to perform the worst among the three types of noise. We aim to demonstrate that our results are robust and superior, not relying on a particular choice of hyper-parameters or experiment scenarios.
Figure 4: Comparison Between Different Noise Methods
@@ D.1 Stage 1: IDDQN Algorithm
13:   Update critics: $\Phi\leftarrow\arg\min_{\Phi}\frac{1}{m}\sum(y-\text{Q}^{DQN}_{\pi^{Q}_{\Phi}}(s,a;\Phi))^{2}$
@@ E.1 Experiment Configurations
Our model was trained using the PyTorch framework [21] on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card. The detailed model configurations are shown as Table 2. During the training phase, the model utilized approximately 8.03 GB of GPU memory. For optimization, we employed the Adam optimizer with an initial learning rate of $10^{-4}$ and a decay rate of 0.99. In Stage 1, the batch size was set to 256, while in Stage 2, it was reduced to 16, due to a sharp decrease in sample amount. Additionally, optimization was performed once every 4 time steps, and in Stage 2, the actor was updated once for every two updates of the critic.
@@ E.2 Simulation Setup
In the simulation, we follow previous works [16; 8; 1], setting the total number of drivers to 1,000, with each car having a capacity of 3 passengers. The maximum waiting time for customers is set to 5 minutes.Each episode lasts 30 minutes, divided into 30 time steps, where each step determines the operations for the subsequent minute. For the TSP route optimization and time estimation, we utilize the OSRM simulator [18], with a default traveling speed of 60 km/h. For solving the bipartite matching, we use the Hungarian algorithm [20], provided in SciPy [30].
@@ E.3 Introduction of Comparative Methods
  • Independent MARL: The DeepPool [1] and BMG-Q [16] utilize a similar IDDQN method as described in Section 3.3.1, with BMG-Q employing GAT [29] to capture the relationships among neighboring agents. Additionally,in the original paper fo DeepPool, the authors used CNN. However, due to differences in the observation space of our task, we replaced it with MLP.
  • Centralized Training Decentralized Execution (CTDE): The HIVES [23] framework introduces a QMIX [23] based method to address the shortcomings of IDDQN, specifically the inadequacy of treating the global Q-value as a simple summation of the individual Q-values of each agent. Enders et al. [8] propose a MASAC [11] based approach, allowing each driver to choose whether to accept an order, thereby preventing low-profit orders from negatively impacting the global income.
  • Centralized Training and Centralized Execution (CTCE): CEVD [3], based on VD [25], innovatively combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, akin to the motivation behind BMG-Q.
@@
Through the experimental results in Fig. 3, we observe that DeepPool [1], serving as one of the earliest benchmarks, demonstrates relatively stable and good performance, suggesting the simplicity and effectiveness of IDDQN features. In contrast, BMG-Q [16] significantly improves performance by utilizing FAT to capture neighboring information. As for HIVES [23] and CEVD [3], while they exhibit relatively good performance in the early stages of training—likely due to their hierarchical structure and centralized training methods—their performance becomes unstable in later stages, with rewards even starting to decrease. This instability may stem from the hierarchical approach not adequately addressing the large network input of the mixture network in QMIX and the lazy agent problem in VD. Additionally, their centralized training approach faces the same data scarcity issues as our method, making convergence more challenging. For Enders et al. [8], we note that their method shows worse performance than others. This may be related to their state processing method during training, where they replace the next state in the replay buffer with the request state from the current state to maintain a consistent agent count across two successive time steps, which appears to be a strong assumption. Finally, for the last three methods, their original papers primarily focus their reward functions on the serving order amount, without incorporating additional terms like ours (which also considers income, outcome, and user satisfaction levels). This makes our scenario more complex and may further reduce the performance of their methods in our setting.

### v2

@@ 3.3.2 Stage 2: Centralized TD3 Training
(i) Actor: In the standard AC framework for discrete action problems, the policy network generates probabilities for each action, from which actions are sampled. However, in the ride-sharing task, this approach encounters two significant challenges: (i) First, the action space is exceedingly large. As shown in Appendix A, a typical scenario with 1,000 couriers and 10 orders can yield nearly $10^{30}$ possible actions. (ii) Second, because the orders vary at each step (including both order volume and content), it is impossible to generate a fixed-dimension action probability vector as is customary in the standard AC framework. (iii) Third, due to the dependency among drivers (an order cannot be assigned to multiple workers concurrently), treating drivers as independent individuals for separate action sampling is impractical.
To address these issues, we impose structural assumptions on the policy function to facilitate its derivation (i.e. the proposed action decomposition strategy). We define $\mathscr{P}_{i,j,t}$ as the probability that worker $i$ chooses order $j$ at time $t$, based on the logit model [logic]. Specifically, we first expand the utility matrix $M_{t}$ generated by the Actor QK-Attention to $\mathcal{M}_{t}=[M_{t},N_{t}]\in\mathbb{R}^{n,m_{t}+1}$, where $N_{t}$ is an $n$-dimensional vector representing the utility of each worker choosing no order. This vector is obtained by processing the output of Actor-BERT with a MLP, expressed as $N_{t}=\text{MLP}([\overline{w}_{1,t},\overline{w}_{2,t},\ldots,\overline{w}_{n,t}])$. This allows us to compute the probability of each worker choosing each action using the logit model, yielding $\mathscr{P}_{t}=\text{Softmax}(\mathcal{M}_{t})$.
Furthermore, we assume that the aggregate policy is derived from the product of these probabilities:
@@
where $\text{z}(\cdot)$ is an increasing function that also depends on the current state $S_{t}$ (which we omit for simplicity), and $\text{h}()$ is defined as $\text{h}(A_{t})=\{(i,j)|a_{i,j,t}=1\}$. This is a reasonable simplification, as it implies that if an action $A_{t}$ has a higher value of $\prod_{i,j\in\text{h}(A_{t})}\mathscr{P}_{i,j,t}$, it has a greater probability of being chosen. We wish to emphasize that the probability $\mathscr{P}$ does not exist in reality but serves as a virtual construct defined by us. We connect the output of the network $\mathcal{M}_{t}$ to the policy $\pi_{\Theta}^{T}$ through this defined probability and the mapping function $\text{z}(\cdot)$. Essentially, we restrict the policy space to a smaller class as defined by Eq. 5 to facilitate optimization and application, as follows.
@@
Optimizing this probability using vanilla TD3 is challenging due to the variable action space and the gap between action probabilities and the selected action (the gradient cannot propagate through them). To address this, we employ an approximate policy gradient optimization method [sutton1999policy]:
@@ 4 Experiment
Method DeepPool [al2019deeppool] BMG-Q [hu2025bmg] HIVES [hao2022hierarchical] Enders et al. [enders2023hybrid] CEVD [bose2023sustainable] Triple-BERT Type Independent CTDE Centralized RL Algorithm IDDQN [van2016deep] IDDQN [van2016deep] QMIX [rashid2020monotonic] MASAC [haarnoja2018soft] VD [^2] The original VD is a CTDE method. However, the CEVD variant modifies it to a centralized version. [sunehag2018value] IDDQN [van2016deep] $\rightarrow$ TD3 [fujimoto2018addressing] Multi-Agent $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark\rightarrow\times$ Reward Type Local Local Global Local Global Local $\rightarrow$ Global Network Backbone MLP [^3] In the original paper fo DeepPool, the authors used CNN. However, due to differences in the observation space of our task, we replaced it with MLP. GAT [velickovic2017graph] GRU [cho2014learning] MLP+Attention MLP BERT [devlin2019bert] Model Size 20K 117K 16M 118K 23K 16M GPU Occupation (GB) 3.97 4.28 6.01 8.19 21.45 8.03 Average Reward ($10^{3}$) 12.72 13.04 12.37 12.04 13.16 14.73
@@
To validate the proposed method, we evaluate its performance in the ride sharing dispatching task using real-world yellow ride-hailing data from Manhattan, New York City [TLCData] [^4] https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page . To illustrate the efficiency and superiority of our proposed Triple-BERT, we compare it with several previous ride sharing methods of different types, including Independent MARL, CTDE MARL, and Centralized MARL, as shown in Table 1. Detailed information regarding our experiment configuration, simulator setup, and a comprehensive description of the comparative experiment can be found in Appendix E.
@@
Figure 3: Training Process: Each method is trained five times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$. The shaded area represents the standard deviation.
@@
As shown in Fig. 3, we first illustrate the training process of different models by evaluating their performance in the training scenario every 10 episodes. The six sub-figures depict the cumulative reward, the number of orders served, and the average delivery time, detour time, pickup time, and confirmation time for each order. It is evident that our method outperforms the other models in most metrics, with the cumulative reward exceeding that of the best alternative method by approximately 15%. The highest number of served orders indicates that our method achieves better cooperation among workers. We then evaluate these methods over different periods, and the average rewards are shown in Table 1, where our method also demonstrates the best performance. More details about the experimental results can be found in Appendix E.4.
To further illustrate the generalization capacity of our method, we evaluate it in additional trials, specifically utilizing all-day data from July 18, 2024, drawn from High Volume For-Hire Vehicle (FHV) trip data [TLCData]. This dataset features a different order distribution compared to the yellow-taxi data used for training. The 24-hour data generates 48 episodes, each lasting 30 minutes, during which the order volume varies from 734 to 5,989 in each episode. The experimental results are presented in Table 2. The findings remain consistent with previous results: our Triple-BERT model achieves the highest reward by optimizing pickup time to improve the service rate, even if this leads to higher delivery and detour times due to increased order bundling. Additionally, we observe that our method exhibits a higher standard deviation in rewards compared to others. This is because, in low-order volume scenarios, performance among different methods does not substantially differ, as all orders can be effectively served. However, in high-order volume scenarios, our Triple-BERT model demonstrates a significant advantage, resulting in higher rewards. Consequently, this contributes to a greater standard deviation when aggregating all scenarios. More expanded experiment could be found at Appendix E.
Table 2: Performance of Different Methods in Manhattan FHV Data [TLCData]
Method Reward Service Rate Delivery Time Detour Time Pickup Time Confirmation Time DeepPool [al2019deeppool] 11258.56±2578.68 0.76±0.18 13.93±0.93 2.54±1.48 10.19±0.78 0.09±0.07 BMG-Q [hu2025bmg] 11899.39±2804.65 0.78±0.17 13.27±1.00 2.44±1.54 9.39±0.43 0.15±0.11 HIVES [hao2022hierarchical] 11183.12±2458.29 0.78±0.18 13.72±1.48 2.87±1.90 9.77±0.68 0.05±0.03 Enders et al. [enders2023hybrid] 10512.65±2744.34 0.78±0.17 14.12±0.48 3.06±0.43 9.92±0.55 1.37±0.99 CEVD [bose2023sustainable] 12556.74±3303.93 0.80±0.13 12.33±0.72 2.25±1.33 8.02±1.27 0.09±0.08 Triple-BERT 14329.74±4627.26 0.88±0.11 13.07±0.61 2.78±0.92 7.02±0.88 0.34±0.32
@@
Then to demonstrate the model’s efficiency, we conduct a series of ablation studies. In terms of model training, we compare the performance of the model with and without stage 1 pre-training. Regarding the network structure, we primarily compare the QK-Attention mechanism with and without the proposed positive normalization module. The detailed results are shown in Fig. 3. We observe that without stage 1 pre-training, the model fails to converge and exhibits significant fluctuations. Particularly in the later stages, the reward begins to decrease, which can be attributed to the lack of samples. Additionally, without the proposed normalization in QK-Attention, the model performs poorly, underperforming compared to all other methods. This is due to parameter redundancy, which leads to substantial fluctuations and hinders efficient learning.
@@ 5 Conclusion
However, compared to traditional MARL-based ride-sharing methods, Triple-BERT is more sensitive to single points of failure, as its decisions depend on comprehensive information from all drivers and orders. Efficient strategies to address this dilemma warrant exploration in future work. Additionally, while this study represents the first centralized SARL-based approach to ride-sharing, we view it as merely the starting point for this new paradigm. Future research could focus on identifying more efficient SARL frameworks or enhancing our existing method, such as exploring importance sampling within our off-policy policy gradient-based actor optimization method, or investigating the use of offline training to replace our pre-training phase.
@@ Appendix C Policy Gradient Proof
According to the policy gradient theory [sutton1999policy], we have:
@@
As introduced in Section 3.3.2, the probability $\mathscr{P}$ and the mapping function $\text{z}(\cdot)$ do not exist in reality but serve as virtual constructs used to connect the output to the policy $\pi_{\Theta}^{T}$ for simplified optimization and application. This means we can define them in any format we choose, and they restrict the policy space to a structural class as defined by Eq. 5.
To further simplify optimization, we define the format of $\text{z}(\cdot)$ as $z(x)=ax^{b}$, where $a,b>0$. This formulation has the advantage that the elasticity of $\text{z}(\cdot)$ is a positive constant, i.e. $\mathcal{E}_{\text{z}(x),x}=ab$. Thus, we have: $\nabla_{\Theta}\text{J}(\Theta)\propto\mathbb{E}_{\pi_{\Theta}^{T}}\left[\left(\text{Q}^{TD3}_{\pi_{\Theta}^{T}}(S_{t},A_{t})-B\right)\nabla_{\Theta}\sum_{i,j\in\text{h}(A_{t})}\log\mathscr{P}_{i,j,t}\right]$, corresponding to Eq. 7. This format of $\text{z}(\cdot)$ aligns with our assumption that $\text{z}(\cdot)$ should be an increasing function, which implies that if a driver-order pair has a higher probability of being chosen, it should also have a higher utility, resulting in a greater likelihood of being selected in the joint action.
@@
As mentioned in Section 3.3.2, during training, we add random noise to $\mathscr{P}_{t}$ and then choose the action that maximizes $\sum_{i,j\in\text{h}(A_{t})}\log\mathscr{P}_{i,j,t}$. Currently, the mapping from $\sum_{i,j\in\text{h}(A_{t})}\log\mathscr{P}_{i,j,t}$ to the choosing probability $\pi_{\Theta}^{T}$ corresponds to $\text{z}(\cdot)$. To further illustrate the robustness of our method, we compare the performance of our model using Gaussian noise [holder2025multi], uniform noise, and binary symmetric channel (BSC) noise, where the noise follows a Bernoulli distribution and has been widely utilized in previous work [hu2025bmg; hu2025coordinating]. During training, we gradually reduce the noise to make the policy more deterministic. The experimental results are shown in Fig. 4, where we observe that, despite certain performance differences between the various types of noise, they all outperform conventional MARL methods. This suggests the efficiency and high robustness of our proposed method, indicating that the detailed expression of $\text{z}(x)$ does not significantly influence the validation of the method based on Eq. 7, even if it may cause some performance gaps. The optimal noise for our task may require further exploration. For fairness, we choose to use BSC noise when comparing with other methods, even though it appears to perform the worst among the three types of noise. We aim to demonstrate that our results are robust and superior, not relying on a particular choice of hyper-parameters or experiment scenarios.
Figure 4: Comparison Between Different Noise Methods: Each method is trained three times, and the curve is smoothed using EMA with $\alpha=0.1$. The shaded area represents the range of fluctuations, while the solid line indicates the average value.
@@ D.1 Stage 1: IDDQN Algorithm
13:   Update Q-Network: $\Phi\leftarrow\arg\min_{\Phi}\frac{1}{m}\sum(y-\text{Q}^{DQN}_{\pi^{Q}_{\Phi}}(s,a;\Phi))^{2}$
@@ E.1 Experiment Configurations
Our model was trained using the PyTorch framework [paszke2019pytorch] on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card. The detailed model configurations are shown as Table 3. During the training phase, the model utilized approximately 8.03 GB of GPU memory. For optimization, we employed the Adam optimizer with an initial learning rate of $10^{-4}$ and a decay rate of 0.99. In Stage 1, the batch size was set to 256, while in Stage 2, it was reduced to 16, due to a sharp decrease in sample amount. Additionally, optimization was performed once every 4 time steps, and in Stage 2, the actor was updated once for every two updates of the critic.
@@ E.2 Simulation Setup
In the simulation, we follow previous works [hu2025bmg; enders2023hybrid; al2019deeppool], setting the total number of drivers to 1,000, with each car having a capacity of 3 passengers. The maximum waiting time for customers is set to 5 minutes.Each episode lasts 30 minutes, divided into 30 time steps, where each step determines the operations for the subsequent minute. For the TSP route optimization and time estimation, we utilize the OSRM simulator [luxen-vetter-2011], with a default traveling speed of 60 km/h. For solving the bipartite matching, we use the Hungarian algorithm [mills2007dynamic], provided in SciPy [virtanen2020scipy].
@@ E.3 Introduction of Comparative Methods
  • Independent MARL: The DeepPool [al2019deeppool] and BMG-Q [hu2025bmg] utilize a similar IDDQN method as described in Section 3.3.1, with BMG-Q employing GAT [velickovic2017graph] to capture the relationships among neighboring agents. Additionally,in the original paper fo DeepPool, the authors used CNN. However, due to differences in the observation space of our task, we replaced it with MLP.
  • Centralized Training Decentralized Execution (CTDE): The HIVES [rashid2020monotonic] framework introduces a QMIX [rashid2020monotonic] based method to address the shortcomings of IDDQN, specifically the inadequacy of treating the global Q-value as a simple summation of the individual Q-values of each agent. Enders et al. [enders2023hybrid] propose a MASAC [haarnoja2018soft] based approach, allowing each driver to choose whether to accept an order, thereby preventing low-profit orders from negatively impacting the global income.
  • Centralized Training and Centralized Execution (CTCE): CEVD [bose2023sustainable], based on VD [sunehag2018value], innovatively combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, akin to the motivation behind BMG-Q.
@@
Through the experimental results in Fig. 3, we observe that DeepPool [al2019deeppool], serving as one of the earliest benchmarks, demonstrates relatively stable and good performance, suggesting the simplicity and effectiveness of IDDQN features. In contrast, BMG-Q [hu2025bmg] significantly improves performance by utilizing FAT to capture neighboring information. As for HIVES [rashid2020monotonic] and CEVD [bose2023sustainable], while they exhibit relatively good performance in the early stages of training—likely due to their hierarchical structure and centralized training methods—their performance becomes unstable in later stages, with rewards even starting to decrease. This instability may stem from the hierarchical approach not adequately addressing the large network input of the mixture network in QMIX and the lazy agent problem in VD. Additionally, their centralized training approach faces the same data scarcity issues as our method, making convergence more challenging. For Enders et al. [enders2023hybrid], we note that their method shows worse performance than others. This may be related to their state processing method during training, where they replace the next state in the replay buffer with the request state from the current state to maintain a consistent agent count across two successive time steps, which appears to be a strong assumption. Finally, for the last three methods, their original papers primarily focus their reward functions on the serving order amount, without incorporating additional terms like ours (which also considers income, outcome, and user satisfaction levels). This makes our scenario more complex and may further reduce the performance of their methods in our setting.